OpenStudy (anonymous):

YAY last one thnks OS! The area of a rectangle can be found using the fromula A=Lw. the area of a square is based on the same concept exept that the sides of a square are equal. use the variable "s" to represent the length of the side of a square. write an area fromula for a square, and solve the equation for "s"

OpenStudy (anonymous):

Formula is A=L*w

OpenStudy (anonymous):

\[s^2\]

OpenStudy (anonymous):

\[s=\sqrt{A}\]

OpenStudy (ghazi):

well just put L=w and you'll have area= l*l=l^2

OpenStudy (ghazi):

\[S=L*L = L^2\]

OpenStudy (anonymous):

now what

OpenStudy (ghazi):

that's it

OpenStudy (anonymous):

A=lw ---> L=w ---> A=l*l=L^2---> S=L*L=L^2

OpenStudy (ghazi):

yes

OpenStudy (phi):

A=L*W that is a good start a square has equal sides so its L and W are equal use the variable "s" to represent the length of the side of a square. so they want you to use s for L and for W: A= s*s or A= s^2 Now they want solve the equation for "s"